OpenStudy (jjuden):

Find the inverse of the function y=3x-8 a y=8-3x b y=3/x+8/3 c y=1/3x-8 d y=x+8/3

OpenStudy (usernamenottaken):

to find the inverse, solve this equation x=3y-8 should help. tell me if it works

OpenStudy (jjuden):

i got y=8/3=x/3 but thats not one of the answers

OpenStudy (jjuden):

plus not =

OpenStudy (usernamenottaken):

huh, i swore that works for me by simply isolating y